• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Solvignon

    Log entry time 00:00:11 on April18,2007

    Entry number 199435

    keyword=Shift summary

    16:00 Start shift with controlled access to fix right Q1.

    16:43 We are now in beam permit

    16:44 Beam in the hall ramping up to 95uA.

    17:18 End-of-run failed. Killed coda. Then restarted. Now no more events are recorded. Killed coda again. Rebooted all components. Now looks OK. (We forgot to do reset between configure and download the fisrt time)

    17:35 Start run# 2588.

    22:08 MCC called. The attenuator is at the max now so they need to
    lower the beam current.

    00:00 Smooth shift.

    -------------------------
    List of runs

    2586 16:51-17:18 98uA, 1.2M End-of run failed
    2588 17:35-18:00 100uA, 1.2M
    2589 18:00-18:23 100uA, 1.2M
    2590 18:23-18:47 100uA, 1.2M
    2591 18:47-19:10 100uA, 1.2M
    2592 19:10-19:36 100uA, 1.2M
    2593 19:36-20:00 100uA, 1.2M
    2594 20:00-20:23 100uA, 1.2M
    2595 20:23-20:49 100uA, 1.2M
    2596 20:49-21:15 100uA, 1.2M
    2597 21:15-21:40 100uA, 1.2M
    2598 21:40-22:06 100uA, 1.2M
    2599 22:07-22:32 95uA, 1.2M
    2600 22:32-23:04 95uA, 1.2M
    2601 23:04-23:31 95uA, 1.2M
    2602 23:31-23:57 95uA, 1.2M
    2603 23:57-...